# Bayesian Multivariable Bidirectional Mendelian Randomization

### v1

@@ 2.3 Priors and Posteriors
The priors are chosen to support both parameter estimation and structural learning, with a particular emphasis on inducing sparsity in the causal graph (via $\mathbf{A}$) and in the confounding structure (via $\mathbf{\Sigma^{*}}$).
Priors on the causal effect matrix $\mathbf{A}$.
We place a spike-and-slab prior on each off-diagonal entry $a_{jh}$ of the matrix $\mathbf{A}$. The presence of a directed edge from $Y_{h}$ to $Y_{j}$ is governed by a binary inclusion variable $\gamma_{jh}$, and the effect size $a_{jh}$ is conditionally modeled as:
$a_{jh}$ $\sim\gamma_{jh}\cdot\text{N}(0,\tau_{jh})+(1-\gamma_{jh})\cdot\text{N}(0,\nu_{1}\cdot\tau_{jh}),$
$\gamma_{jh}$ $\sim\text{Bernoulli}(\rho_{jh})$
$\rho_{jh}$ $\sim\text{Beta}(a_{\rho},b_{\rho})$
$\sqrt{\tau_{jh}}$ $\sim\mathcal{C}^{+}(0,1),$
where $\mathcal{C}^{+}(0,1)$ is the standard half-Cauchy distribution. Here, we fix $\nu_{1}\ll 1$ to ensure good separation between signals and noises. Following makalic2016simple , half-Cauchy distribution can be reparameterized as:
$x\sim\mathcal{C}^{+}(0,1)\implies x^{2}\mid a\sim\text{IG}(1/2,1/a),\quad a\sim\text{IG}(1/2,1),$
where $\text{IG}(\cdot,\cdot)$ denotes the inverse-gamma distribution. This allows a closed-form Gibbs update.
Priors on the instrument effect matrix $\mathbf{B}$.
When there may be horizontal pleiotropy (the violation of the exclusion restriction assumption) or weak instruments (the violation of the relevance assumption), we do not impose fixed structural zeros in $\mathbf{B}$ but instead adopt a spike-and-slab prior for each entry of $\mathbf{B}$:
$b_{jh}$ $\sim\phi_{jh}\cdot\text{N}(0,\eta_{jh})+(1-\phi_{jh})\cdot\text{N}(0,\nu_{2}\cdot\eta_{jh}),$
$\phi_{jh}$ $\sim\text{Bernoulli}(\psi_{jh}),$
$\psi_{jh}$ $\sim\text{Beta}(a_{\psi},b_{\psi}),$
$\sqrt{\eta_{jh}}$ $\sim\mathcal{C}^{+}(0,1),$
where $\nu_{2}\ll 1$ and $\phi_{jh}$ indicates whether $X_{h}$ is a (significant) instrument for $Y_{j}$. This prior formulation encourages sparsity in $\mathbf{B}$, facilitating automatic selection of valid and relevant instruments. Under the InSIDE assumption bowden2015mendelian ; yuan2020testing that the instrument effects are independent of each other, we obtain unbiased causal effect estimation. We call MR.RGM with selection on instruments MR.RGM+. Otherwise, simple normal priors are placed on the non-zero entries of $\mathbf{B}$.
Prior on the covariate effect matrix $\mathbf{C}$.
We assume a conjugate matrix-normal prior on $\mathbf{C}\in\mathbb{R}^{p\times l}$:
$\mathbf{C}\sim\text{MN}_{p\times l}(\mathbf{0},\,\mathbf{\Sigma}^{*},\,\tau\mathbf{I}_{l}).$
Prior on the error covariance matrix $\mathbf{\Sigma^{*}}$.
The error covariance matrix $\mathbf{\Sigma}^{*}=\mathbf{D}\mathbf{D}^{T}+\mathbf{\Sigma}$ combines unmeasured confounding (via $\mathbf{D}$) and measurement noise (via $\mathbf{\Sigma}$). Since we do not know a priori the number $t$ of unmeasured confounders, we directly model $\mathbf{\Sigma}^{*}$ instead of modeling $\mathbf{D}$ and $\mathbf{\Sigma}$. We introduce a binary indicator matrix $\mathbf{Z}=(z_{jh})\in\{0,1\}^{p\times p}$ that encodes whether the off-diagonal entry $\sigma^{*}_{jh}$ is nonzero. The priors are specified as follows:
$\sigma^{*}_{jh}\sim z_{jh}\cdot\text{N}(0,\omega_{1}^{2})+(1-z_{jh})\cdot\text{N}(0,\omega_{2}^{2}),\quad j<h,$
$z_{jh}\sim\text{Bernoulli}(\pi),$
$\sigma^{*}_{jj}\sim\text{Exp}\left(\frac{\lambda}{2}\right),$
subject to the positive-definiteness of the resulting $\mathbf{\Sigma^{*}}$. The hyperparameters $\omega_{1},\omega_{2},\pi$ and $\lambda$ control the sparsity and scale of the inferred error covariance. Following wang2015 , we recommend setting $\omega_{2}\geq 0.01$ and $\omega_{1}/\omega_{2}\leq 1000$ to ensure numerical stability and encourage separation between signals and noises. The parameter $\pi$ lies between 0 and 1, with smaller values favoring sparser structures. For the exponential prior, $\lambda$ may be set to moderately large values such as 5 or 10.
Posterior Computation.
@@
We use MCMC to draw posterior samples of the model parameters. Step-by-step updating scheme is detailed in the Supplementary Materials. Our implementation leverages Rcpp and linear algebra for efficient matrix computations.
@@ 3 Simulation Studies
In this section, we evaluate the performance of the proposed MR.RGM, using simulations designed to reflect complex causal structures commonly found in biological systems. We compare MR.RGM with several baseline methods such as MR packages OneSampleMR, mrbayes, and MendelianRandomization, which includes MR approaches based on Simple Median, Weighted Median and IVW. These baselines offer a diverse representation of current MR tools; however, none explicitly model feedback loops. For clarity, we denote the Simple Median, Weighted Median, and IVW methods from the MendelianRandomization package as MR-SimpleMedian, MR-WeightedMedian, and MR-IVW, respectively. For ablation, we also consider two variants of MR.RGM, namely, MR.RGM_NoConf, which assumes no latent confounding and was implemented by sarkar2025mrrgm , and MR.RGM+, which accounts for horizontal pleiotropy. For the implementation of MR.RGM+, rather than prespecifying instrument-trait mapping, we treat every SNP as a potential instrument for every gene/trait.
@@
Given the graph and IV structure, we simulate data from (15) without observed covariates, where the non-zero off-diagonal entries of $\mathbf{A}$ are sampled from $\text{Uniform}[-0.1,0.1]$, reflecting small effect sizes to capture low-signal conditions common in real data, the non-zero entries of $\mathbf{B}$ are set to be 1, the entries of $\mathbf{D}$ are sampled from $\{-1,+1\}$, the instruments and the unmeasured confounders are drawn from independent standard normal distributions, and the errors are drawn from independent centered normal distributions with variance 9.
For Case III, since the competing methods do not account for horizontal pleiotropy, we randomly assign each instrument to one of the two traits if an instrument affects two traits.
We vary sample size $n\in\{500,1000,10000,30000\}$ and number of traits $p\in\{5,10\}$, yielding 8 settings per graph structure. Each simulation is repeated 20 times. Performance is evaluated based on the following criteria:
@@
  • Confounding Structure Recovery: We normalize the true variance-covariance matrix $\mathbf{\Sigma}^{*}=\mathbf{DD}^{T}+\mathbf{\Sigma}$ to the range of $[0,1]$ and threshold it at its empirical mean to create a true confounding structure. We then compute AUC, TPR, FDR, and MCC of our estimated confounding structure.
  • Instrument-Trait Selection Accuracy. When MR.RGM+ is applied (i.e., in Case III), we also report the AUC of the instrument-trait selection.
In the main text, we focus on visualizing the AUC for graph recovery, confounding structure recovery, and instrument-trait selection, and the MAD for causal effect estimation, for network size p=10p{=}10. The corresponding plots for network size p=5p{=}5 as well as the full set of performance tables for all metrics and scenarios are provided in the Supplementary Materials.
Results for Cases I-II.
The results for Cases I and II are similar and hence are reported together. Figures  1(a) and 2(a) show that MR.RGM consistently achieves the highest AUC in graph recovery across all sample sizes in both cases. MR.RGM_NoConf underperforms MR.RGM for large $n$, reflecting the influence of unmeasured confounding. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR and mrbayes, improve steadily with increasing $n$ and nearly catch up MR.RGM by n=10,000n{=}10{,}000-30,00030{,}000.
Figures 1(b) and 2(b) show that MR.RGM attains the lowest MAD for the causal effect estimation for all sample sizes. MR.RGM_NoConf is competitive at $n\in\{500,1000\}$ but become less so as $n$ grows. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR, and mrbayes improve with sample size and eventually surpass MR.RGM_NoConf at larger $n$, while remaining short of MR.RGM.
@@
Figures 3(a) and 3(b) show that MR.RGM recovers the confounding structure increasingly well with sample size: the median AUC approaches 1.0 for $n\geq 10{,}000$. Competing methods do not infer the confounding structure.
@@
Scalability Analysis.
We benchmark the runtime of MR.RGM against the competing methods. We fix the number of observations at the largest value used in our simulations, n=30,000n=30{,}000, and vary the number of traits $p\in\{2,5,10,20\}$. All Bayesian methods (MR.RGM, MR.RGM_NoConf, MR.RGM+, mrbayes) use 50,000 MCMC iterations with 10,000 burn-in; for MendelianRandomization we also run 50,000 iterations. Benchmarks were executed in RStudio on an Apple M2 Pro machine (10-core CPU, 3.5 GHz) with 16 GB unified memory. Each method is run 20 times, and we report the median wall-clock runtime in seconds in Figure 7. As $p$ increases, the runtime rises for all methods, but the MR.RGM family remains practical. For example, at $p=20$, the median runtime is approximately $90$s for MR.RGM, $55$s for MR.RGM_NoConf, $233$s for MR.RGM+.
Figure 7: Median runtime versus number of traits. Lines show median wall-clock runtime (seconds) over 20 runs for each method as $p$ increases ($p\in\{2,5,10,20\}$) with fixed n=30,000n=30{,}000. Experiments were conducted in RStudio on an Apple M2 Pro machine (10-core CPU, 3.5 GHz) with 16 GB unified memory.

### v2

@@ 2.3 Priors and Posteriors
The priors are chosen to support both parameter estimation and structural learning, with a particular emphasis on inducing sparsity in the causal graph ($\mathbf{A}$) and in the confounding structure ($\mathbf{\Sigma^{*}}$) via two sets of spike-and-slab priors. A full specification of all prior distributions and hyperparameters is provided in Supplementary Section-S22. As an extension to account for horizontal pleiotropy and weak instruments, we impose another set of spike-and-slab priors to select valid instruments, for which the details are also provided in Supplementary Section-S22.
@@
We use MCMC to draw posterior samples of the model parameters. Step-by-step updating scheme is detailed in the Supplementary Section-S33. Our implementation leverages Rcpp and linear algebra for efficient matrix computations. Figure 2 provides a schematic overview of the full MR.RGM workflow.
[…]
Figure 2: Overview of the proposed MR.RGM workflow. The method can take either individual-level data $(\mathbf{Y,X,U})$ or summary-level second-moment statistics plus sample size $n$ as input. These are plugged into a structural equation model with directed cycles and correlated errors. Spike-and-slab priors and $MC$ are then used to infer the causal graph, the latent confounding structure, and the causal effects, along with their corresponding uncertainties.
2.4 Identifiability of MR.RGM+ Under Instrument Diversity and Sparsity
When there is horizontal pleiotropy (i.e., the violation of the exclusion restriction condition), the proposed model is not identifiable without additional assumptions. Without loss of generality, we drop the observed covariates $\mathbf{U}$ from the proposed SEM:
$\mathbf{Y}=\mathbf{A}\mathbf{Y}+\mathbf{B}\mathbf{X}+\mathbf{E}^{*},\qquad\mathbf{E}^{*}\sim N_{p}(\mathbf{0},\boldsymbol{\Sigma}^{*}),$ (18)
where $\mathrm{diag}(\mathbf{A})=\mathbf{0}$ and $\mathbf{I}_{p}-\mathbf{A}$ is assumed to be invertible. Define the reduced-form coefficient
$\boldsymbol{\Pi}:=(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{B}.$ (19)
By theory of multivariate linear regression, $\boldsymbol{\Pi}$ is identifiable from second moments, $\boldsymbol{\Pi}=\mathrm{Cov}(\mathbf{Y},\mathbf{X})\mathrm{Var}(\mathbf{X})^{-1}$ provided that $\mathrm{Var}(\mathbf{X})$ is invertible. To establish identifiability of $(\mathbf{A},\mathbf{B})$ from $\boldsymbol{\Pi}$, we make the following three assumptions:
Assumption 1 (Instrument Diversity).
All $(p-1)\times(p-1)$ minors of $\boldsymbol{\Pi}$ are non-zero. That is, for every subset $R\subseteq\{1,\ldots,p\}$ with |R|=p1|R|=p-1 and every subset $H\subseteq\{1,\ldots,k\}$ with |H|=p1|H|=p-1:
$\det(\boldsymbol{\Pi}_{R,H})\neq 0.$
We provide a few examples in Supplementary Section-S44. Intuitively, this assumption requires that for each trait $j$, the instruments must provide “diverse” information about the other traits; in other words, no (p1)(p-1) instruments can have collinear effects on the non-$j$ traits. Note that the Instrument Diversity assumption is a condition on $\boldsymbol{\Pi}=(\mathbf{I}-\mathbf{A})^{-1}\mathbf{B}$, whereas the Relevance assumption of a classic MR is a condition on $\mathbf{B}$. Even if row $j$ of $\mathbf{B}$ is zero (no instruments directly affect trait $j$), row $j$ of $\boldsymbol{\Pi}$ can be non-zero through mediation:
$\boldsymbol{\Pi}_{j,:}=\sum_{h}[(\mathbf{I}-\mathbf{A})^{-1}]_{jh}\mathbf{B}_{:,h},$
where $\mathbf{B}_{:,h}$ denotes the $h$th column of $\mathbf{B}$. Hence, a trait with no direct instruments can still have non-zero reduced-form effects. Let $\|\cdot\|_{0}$ denote the $L_{0}$ norm, i.e., the number of non-zero elements in a vector, and let $\mathbf{B}_{j,:}$ denote the $j$th row of $\mathbf{B}$.
Assumption 2 (Sparsity).
Every trait has at most $\frac{k-p+1}{2}$ instruments, i.e., $\max_{j\in\{1,\ldots,p\}}\|\mathbf{B}_{j,:}\|_{0}\leq\frac{k-p+1}{2}$.
Intuitively, this assumption requires the violation of the Exclusion Restriction assumption to be sparse. In other words, a dense $\mathbf{B}$ implies a lot of violations, which is ruled out by the Sparsity assumption. For example, for $p=10$ traits and $k=30$ instruments, we assume $\|\mathbf{B}_{j,:}\|_{0}\leq\frac{30-10+1}{2}=10.5$, i.e., each trait can have at most $10$ instruments with direct effects.
Assumption 3 (Independence).
The instruments are not confounded.
This assumption is also required in the classic MR.
Theorem 1 (Identifiability of MR.RGM+).
Under Assumptions 1-3, $(\mathbf{A},\mathbf{B},\boldsymbol{\Sigma}^{*})$ is identifiable.
The proof is provided in Supplementary Section-S4; here we provide a sketch. The reduced-form matrix $\boldsymbol{\Pi}=(\mathbf{I}_{p}-\mathbf{A})^{-1}\mathbf{B}$ is identified from data, but without constraints, infinitely many $(\mathbf{A},\mathbf{B})$ pairs yield the same $\boldsymbol{\Pi}$. The key insight is that the sparsity constraint on $\mathbf{B}$ provides enough equations to identify $\mathbf{A}$. Inverting the relationship gives $\mathbf{B}=(\mathbf{I}_{p}-\mathbf{A})\boldsymbol{\Pi}$, so each entry satisfies:
$b_{jh}=\pi_{jh}-\sum_{i\neq j}a_{ji}\pi_{ih}.$
When $b_{jh}=0$ (instrument $h$ has no direct effect on trait $j$), this becomes a linear constraint on the causal effects $\{a_{ji}\}_{i\neq j}$:
$\pi_{jh}=\sum_{i\neq j}a_{ji}\pi_{ih}.$
Crucially, the constraints for trait $j$ involve only the p1p-1 unknowns in row $j$ of $\mathbf{A}$, so the problem decomposes into $p$ independent linear systems. For trait $j$ with $s_{j}:=\|\mathbf{B}_{j,:}\|_{0}$ direct instrument effects, there are $k-s_{j}$ zero constraints in p1p-1 unknowns. With $s_{j}\leq(k-p+1)/2$, any two candidate solutions must share at least $k-2s_{j}\geq p-1$ zero constraints. This overdetermined system can have at most one solution. The Instrument Diversity condition ensures these shared constraints have full rank, guaranteeing that the unique solution exists.
@@ 3 Simulation Studies
We evaluate the performance of the proposed MR.RGM, using simulations designed to reflect complex causal structures commonly found in biological systems. We compare MR.RGM with several baseline methods such as MR packages OneSampleMR, mrbayes, and MendelianRandomization, which include MR approaches based on Simple Median, Weighted Median, and IVW. These baselines offer a diverse representation of current MR tools; however, none explicitly model feedback loops. For clarity, we denote the Simple Median, Weighted Median, and IVW methods from the MendelianRandomization package as MR-SimpleMedian, MR-WeightedMedian, and MR-IVW, respectively. For ablation, we also consider two variants of MR.RGM, namely, MR.RGM_NoConf, which assumes no latent confounding and was implemented by Sarkar and Ni (2025), and MR.RGM+, which accounts for horizontal pleiotropy. For the implementation of MR.RGM+, rather than prespecifying instrument-trait mapping, we treat every SNP as a potential instrument for every gene/trait.
@@
Given the graph and IV structure, we simulate data from (15) without covariates, where the non-zero off-diagonal entries of $\mathbf{A}$ are sampled from $\text{Uniform}[-0.1,0.1]$, reflecting small effect sizes to capture low-signal conditions common in real data. To ensure that both the instrument effects and the confounding effects remain detectable relative to these weak causal signals, the non-zero entries of $\mathbf{B}$ are set to be $1$, the entries of $\mathbf{D}$ are sampled from $\{-1,+1\}$, giving them a larger scale than the entries of $\mathbf{A}$. The instruments and the unmeasured confounders are drawn from independent standard normal distributions. The errors are drawn from independent centered normal distributions with variance $9$, resulting in a weak signal-to-noise setting.
For Case $3$, since the competing methods do not account for horizontal pleiotropy, we randomly assign each instrument to one of the two traits if an instrument affects two traits.
We vary sample size $n\in\{500,1000,10000,30000\}$ and number of traits $p\in\{5,10\}$, yielding $8$ settings per scenario. Each simulation is repeated $20$ times. Performance is evaluated based on the following criteria:
@@
  • Confounding Structure Recovery: We normalize the true variance-covariance matrix $\mathbf{\Sigma}^{*}=\mathbf{DD}^{T}+\mathbf{\Sigma}$ to the range of $[0,1]$ and threshold it at its empirical mean to create a “true” significant confounding structure. We then compute AUC, TPR, FDR, and MCC of our estimated confounding structure.
  • Instrument-Trait Selection Accuracy. When MR.RGM+ is applied (i.e., in Case $3$), we also report the AUC of the instrument-trait selection.
In the main text, we focus on visualizing the AUC for graph recovery, confounding structure recovery, and instrument–trait selection, and the MAD for causal effect estimation, for network size p=10p{=}10. In addition, for the scale-free network with feedback loops and unmeasured confounding and network size p=10p{=}10, we report the full performance tables for graph recovery (AUC, TPR, FDR, and MCC; Table 1) and for causal effect estimation (MaxAbsDev, MeanAbsDev, and MeanSqDev; Table 2). The corresponding plots for network size p=5p{=}5, as well as the remaining performance tables for both p=5p{=}5 and p=10p{=}10 across all scenarios, are provided in Supplementary Section-S55.
Results for Cases 11-22.
The results for Cases $1$ and $2$ are similar and hence are reported together. Figures 3(a) and 4(a) show that MR.RGM consistently achieves the highest AUC in graph recovery across all sample sizes in both cases. MR.RGM_NoConf underperforms MR.RGM for large $n$, reflecting the influence of unmeasured confounding. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR and mrbayes, improve steadily with increasing $n$ and nearly catch up MR.RGM by n=10,000n{=}10{,}000-30,00030{,}000.
Figures 3(b) and 4(b) show that MR.RGM attains the lowest MAD for the causal effect estimation for all sample sizes. MR.RGM_NoConf is competitive at $n\in\{500,1000\}$ but become less so as $n$ grows. MR-SimpleMedian, MR-WeightedMedian, MR-IVW, OneSampleMR, and mrbayes improve with sample size and eventually surpass MR.RGM_NoConf at larger $n$, while remaining short of MR.RGM.
Tables 1 and 2 summarize the numerical results for Case $1$ with p=10p{=}10, and they closely mirror the patterns seen in the AUC and MAD boxplots. Table 1 shows that MR.RGM achieves comparable or higher TPR together with consistently lower FDR and higher MCC than competing methods as the sample size increases. Table 2 likewise shows that MR.RGM attains smaller MaxAbsDev and MeanSqDev than all alternatives. By contrast, MR.RGM_NoConf lags behind in both graph recovery and causal effect estimation, especially at larger $n$, reflecting its inability to account for unmeasured confounding.
@@
Setting Method AUC TPR FDR MCC
n = $500$ MR.RGM 0.732 (0.055) 0.472 (0.108) 0.536 (0.096) 0.331 (0.114)
MR.RGM_NoConf 0.718 (0.045) 0.494 (0.085) 0.605 (0.058) 0.281 (0.077)
MR-SimpleMedian 0.612 (0.047) 0.093 (0.067) 0.397 (0.264) 0.161 (0.109)
MR-WeightedMedian 0.601 (0.073) 0.111 (0.072) 0.500 (0.278) 0.154 (0.136)
MR-IVW 0.631 (0.057) 0.182 (0.092) 0.508 (0.219) 0.199 (0.135)
OneSampleMR 0.633 (0.060) 0.188 (0.093) 0.506 (0.223) 0.205 (0.141)
mrbayes 0.630 (0.058) 0.185 (0.093) 0.504 (0.218) 0.203 (0.134)
n = 1000 MR.RGM 0.824 (0.067) 0.537 (0.096) 0.428 (0.100) 0.445 (0.111)
MR.RGM_NoConf 0.789 (0.073) 0.585 (0.110) 0.547 (0.090) 0.371 (0.121)
MR-SimpleMedian 0.681 (0.050) 0.164 (0.075) 0.358 (0.206) 0.246 (0.107)
MR-WeightedMedian 0.695 (0.057) 0.188 (0.079) 0.402 (0.184) 0.251 (0.107)
MR-IVW 0.731 (0.055) 0.265 (0.097) 0.454 (0.203) 0.281 (0.152)
OneSampleMR 0.732 (0.055) 0.259 (0.087) 0.439 (0.212) 0.284 (0.146)
mrbayes 0.731 (0.055) 0.265 (0.097) 0.463 (0.201) 0.277 (0.151)
n = 10000 MR.RGM 0.993 (0.017) 0.972 (0.042) 0.033 (0.041) 0.962 (0.046)
MR.RGM_NoConf 0.918 (0.030) 0.906 (0.080) 0.485 (0.060) 0.579 (0.072)
MR-SimpleMedian 0.992 (0.007) 0.957 (0.047) 0.139 (0.082) 0.882 (0.071)
MR-WeightedMedian 0.994 (0.005) 0.969 (0.042) 0.128 (0.083) 0.897 (0.066)
MR-IVW 0.998 (0.003) 0.994 (0.017) 0.175 (0.078) 0.878 (0.054)
OneSampleMR 0.998 (0.003) 0.994 (0.017) 0.172 (0.072) 0.881 (0.050)
mrbayes 0.998 (0.003) 0.994 (0.017) 0.174 (0.076) 0.880 (0.053)
n = 30000 MR.RGM 0.994 (0.017) 0.988 (0.030) 0.000 (0.000) 0.992 (0.019)
MR.RGM_NoConf 0.939 (0.030) 0.920 (0.052) 0.452 (0.056) 0.617 (0.058)
MR-SimpleMedian 1.000 (0.000) 1.000 (0.000) 0.140 (0.068) 0.907 (0.047)
MR-WeightedMedian 1.000 (0.000) 1.000 (0.000) 0.154 (0.077) 0.897 (0.054)
MR-IVW 1.000 (0.000) 1.000 (0.000) 0.201 (0.061) 0.864 (0.044)
OneSampleMR 1.000 (0.000) 1.000 (0.000) 0.201 (0.061) 0.864 (0.044)
mrbayes 1.000 (0.000) 1.000 (0.000) 0.200 (0.060) 0.865 (0.043)
Table 1: Graph recovery performance in a scale-free network with feedback loops and unmeasured confounding, with network size $p=10$.
Setting Method MaxAbsDev MeanAbsDev MeanSqDev
n = 500 MR.RGM 0.173 (0.043) 0.035 (0.004) 0.002 (0.001)
MR.RGM_NoConf 0.179 (0.045) 0.039 (0.003) 0.003 (0.001)
MR-SimpleMedian 0.315 (0.053) 0.094 (0.009) 0.014 (0.002)
MR-WeightedMedian 0.309 (0.059) 0.090 (0.009) 0.013 (0.002)
MR-IVW 0.287 (0.051) 0.081 (0.006) 0.010 (0.002)
OneSampleMR 0.285 (0.049) 0.081 (0.006) 0.010 (0.002)
mrbayes 0.286 (0.049) 0.081 (0.006) 0.010 (0.002)
n = 1000 MR.RGM 0.130 (0.018) 0.030 (0.002) 0.002 (0.0003)
MR.RGM_NoConf 0.154 (0.023) 0.035 (0.002) 0.002 (0.0002)
MR-SimpleMedian 0.225 (0.037) 0.064 (0.006) 0.007 (0.001)
MR-WeightedMedian 0.215 (0.037) 0.061 (0.005) 0.006 (0.001)
MR-IVW 0.191 (0.033) 0.056 (0.004) 0.005 (0.001)
OneSampleMR 0.192 (0.031) 0.056 (0.004) 0.005 (0.001)
mrbayes 0.192 (0.032) 0.056 (0.004) 0.005 (0.001)
n = 10000 MR.RGM 0.059 (0.015) 0.011 (0.001) 0.0003 (0.0001)
MR.RGM_NoConf 0.102 (0.013) 0.029 (0.002) 0.001 (0.0001)
MR-SimpleMedian 0.071 (0.011) 0.021 (0.002) 0.001 (0.0001)
MR-WeightedMedian 0.069 (0.011) 0.020 (0.002) 0.001 (0.0001)
MR-IVW 0.063 (0.009) 0.018 (0.002) 0.001 (0.0001)
OneSampleMR 0.063 (0.010) 0.018 (0.002) 0.0005 (0.0001)
mrbayes 0.063 (0.010) 0.018 (0.002) 0.0005 (0.0001)
n = 30000 MR.RGM 0.030 (0.018) 0.006 (0.001) 0.00007 (0.00003)
MR.RGM_NoConf 0.094 (0.011) 0.028 (0.002) 0.001 (0.0001)
MR-SimpleMedian 0.041 (0.005) 0.013 (0.001) 0.0002 (0.0001)
MR-WeightedMedian 0.040 (0.005) 0.013 (0.001) 0.0002 (0.0001)
MR-IVW 0.038 (0.006) 0.011 (0.001) 0.0002 (0.0001)
OneSampleMR 0.037 (0.005) 0.011 (0.001) 0.0002 (0.00003)
mrbayes 0.037 (0.005) 0.011 (0.001) 0.0002 (0.00003)
Table 2: Causal effect estimation performance in a scale-free network with feedback loops and unmeasured confounding, with network size $p=10$.
Figures 5(a) and 5(b) show that MR.RGM recovers the confounding structure increasingly well with sample size: the median AUC approaches $1.0$ for $n\geq 10{,}000$. Competing methods do not infer the confounding structure.
@@
Scalability Test.
We benchmark the runtime of MR.RGM against the competing methods. We fix the number of observations at the largest value used in our simulations, n=30,000n=30{,}000, and vary the number of traits $p\in\{2,5,10,20\}$. All Bayesian methods (MR.RGM, MR.RGM_NoConf, MR.RGM+, mrbayes) use 50,00050{,}000 MCMC iterations with 10,00010{,}000 burn-in; for MendelianRandomization we also run 50,00050{,}000 iterations. Benchmarks were executed in RStudio on an Apple M22 Pro machine ($10$-core CPU, $3.5$ GHz) with $16$ GB unified memory. Each method is run $20$ times, and we report the median wall-clock runtime in seconds in Figure 9. As $p$ increases, the runtime rises for all methods, but the MR.RGM family remains practical. For example, at $p=20$, the median runtime is approximately $90$s for MR.RGM, $55$s for MR.RGM_NoConf, $233$s for MR.RGM+.
Figure 9: Median runtime versus number of traits. Lines show median wall-clock runtime (seconds) over 2020 runs for each method as $p$ increases ($p\in\{2,5,10,20\}$) with fixed n=30,000n=30{,}000. Experiments were conducted in RStudio on an Apple M22 Pro machine ($10$-core CPU, $3.5$ GHz) with $16$ GB unified memory.